Solve the system by addition
5x-3y= 13
4x-3y=11

Answer by uma(370)


5x-3y= 13
4x-3y=11
Subtracting both the equations,
5x - 4x -3y + 3y = 13 - 11
==> x = 2
Substituting this value of x in the first equation,
5(2) - 3y = 13
==> 10 - 3y = 13
==> -3y = 13 - 10
==> - 3y = 3
==> y = 3/-3
==> y = -1
Thus x = 2 and y = -1 is the solution.